OpenStudy (anonymous):

solve for r p=4r+3t which property is illustrated by 2x^2(x+5)=2x^3+10x^2

OpenStudy (anonymous):

Distributive I think, but I was never good at memorizing names of techniques just teh techniques themselves

OpenStudy (anonymous):

Like ChmE said, it's the distributive property of addition (multiplication distributes over addition).

OpenStudy (anonymous):

ok what about the first one

OpenStudy (anonymous):

p=4r+3t p-3t=4r r = (p-3t)/4 \[r = \frac{p-3t}{4}\]

OpenStudy (anonymous):

thanksss